OpenStudy (inkyvoyd):

Representing the square root of a number n as an infinite fraction. Alright, I want to understand how I can use solving quadratics as continued fractions to approximate the square root of a number.

OpenStudy (inkyvoyd):

if x^2-n=0, x=n/x using recursion, x=n/(n/x) x=n/(n/(n/x)) How exactly would I make an approximation?

OpenStudy (inkyvoyd):

More info at wikipedia's page on solving quadratics with continued fractions.

OpenStudy (anonymous):

maybe like this: \[x^2 - n = 0\]\[x^2 = n\] \[x = \frac{n}{x}\] \[2x = \frac{n}{x} + x \] \[x= \frac{(n/x) + x}{2}\] \[x_{n+1}= \frac{(n/x_{n}) + x_{n}}{2} \]

OpenStudy (anonymous):

pretty proud of that, just came up with it on the spot :D

OpenStudy (inkyvoyd):

OMG THATS THE BABYLONIAN METHOD

OpenStudy (anonymous):

so eg n = 2 start with \[x_0 = 1\] x is approximately: (2/(((2/1)+1)/2))) + ((2/1)+1)/2)/2 (Bracket overload)

OpenStudy (anonymous):

babylonian? cool!

OpenStudy (anonymous):

i just tried making a quadratic with x as a root

OpenStudy (anonymous):

is it important?